OpenStudy (anonymous):

if y varies inversely as x, and y=4 as x=8 find y for the x-value of 2.

OpenStudy (anonymous):

answers: 16 10 8 14

OpenStudy (anonymous):

Do you remember what the formula for an inverse relation is?

OpenStudy (anonymous):

no, i just havent done any of these yet

OpenStudy (anonymous):

Remember if y varies inversely with x, y=k/x. If y varies directly with x, y=kx.

OpenStudy (anonymous):

Ingrain those into your head! Now that you have the formula, plug in x and y to find what k is.

OpenStudy (anonymous):

im not sure how to plug this one in..

OpenStudy (anonymous):

You have y=4 and x=8 correct? You also have y=k/x. We want to know what k is. 4=k/8. k=32. k is a CONSTANT, meaning it doesn't change. Think of it as the opposite to Kim Kardashian's marriages, it lasts forever. Now that we have k=32, the question asks us to figure out what y is when x=2. What do you think y is?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

am i right?

OpenStudy (solomonzelman):

if y=4 and x=8 plug this into y=kx, find the k, and then solve for y=kx when x is 2

OpenStudy (anonymous):

4=32/8?

OpenStudy (anonymous):

did i plug that in correct?

OpenStudy (solomonzelman):

y=4 and x=8 plug this into y=kx so 4=k8 therefore k is 1/2 now find the y when x is 2.

OpenStudy (anonymous):

idk. i got the answer as 16

OpenStudy (solomonzelman):

for y?

OpenStudy (solomonzelman):

y=1/2x or so 1/2(2) =?

OpenStudy (solomonzelman):

@meg_72, do you see why it's not 16?

OpenStudy (anonymous):

the answer i get from that is not listed on my paper

OpenStudy (solomonzelman):

you got 1, right? of your answer is not on the list, then either the question is typed the wrong way, or I am just CRAZY, I think the first....

OpenStudy (anonymous):

lol the answers they gave me are 16 10 8 14

OpenStudy (solomonzelman):

But this is impossible, I am betting any money, on the table, it is 1!

OpenStudy (solomonzelman):

look, 4=k8 so k=1/2 2(1/2) is what?

OpenStudy (solomonzelman):

Do you understand why I keep getting 1?

OpenStudy (anonymous):

haha yea ima bout to find out the answer

OpenStudy (solomonzelman):

$ ∞ it is 1

OpenStudy (anonymous):

it says 16 lol

OpenStudy (solomonzelman):

@wio, @dan815

OpenStudy (solomonzelman):

@meg_72, do you believe that is right?

OpenStudy (solomonzelman):

For real?!

OpenStudy (anonymous):

so im guessing 4=k/8 k=32 32/8 gives 4 so 4=4 then gives us 16 lol idk

OpenStudy (solomonzelman):

k=32? How in the world do you get that?

OpenStudy (solomonzelman):

OH, k/8 not k8, my bad!

OpenStudy (solomonzelman):

you are right, I suck!!!!!!!!!!!!!!